OpenStudy (anonymous):

3x+y=6 4x+2y=6 Use substitution method

OpenStudy (caozeyuan):

so y=6-3x

OpenStudy (caozeyuan):

so 4x+2(6-3x)=6

OpenStudy (caozeyuan):

4x+12-6x=6,2x=6,x=3

OpenStudy (caozeyuan):

y=6-3*3=-3

OpenStudy (uri):

3x+y=6 (-2) 4x+2y=6 ====== -6x-2y=-12 4x+2y=12 ======== -2x = 0 No solution.